# What Breaks Knowledge Graph based RAG? Empirical Insights into Reasoning under Incomplete Knowledge

### v1

@@ 5.4 Influence of Entity Labeling
To assess how different entity labeling schemes affect KG-RAG performance, we evaluate three settings for representing entities in the input: (1) Private ID—randomly assigned IDs with no semantic content, (2) Entity ID—official Freebase IDs (e.g., /m/02mjmr for Barack Obama), and (3) Text Label—natural language labels of entities. Figure 4 shows the F1-scores of all models under each representation, for both incomplete (top) and complete (bottom) KG settings. We observe two key trends: First, text labels significantly boost performance. Models consistently achieve higher scores when entity labels are expressed in natural language. This suggests that LLMs can effectively leverage their internal knowledge when text labels are provided. Second, entity IDs provide limited benefit over random IDs. Surprisingly, using official entity IDs like /m/02mjmr results in performance nearly identical to that of randomly assigned private IDs. This indicates that, despite LLMs potentially memorizing mappings between surface forms and IDs, they are unable to reason with these identifiers in generation tasks. Instead, they appear to treat IDs as opaque tokens unless the text label is explicit.
Figure 4: F1-Score comparison of models using Private ID, Entity ID, and Text Label representations on FB15K-237 benchmark. (a) Incomplete KG. (b) Complete KG.
@@ 5.5 Case Study
To gain deeper insight into the limitations of current KG-RAG methods, we conduct a qualitative analysis of representative failure cases from our benchmark. Table 4 presents two illustrative examples, each highlighting a distinct failure pattern: (1) failure to retrieve relevant reasoning paths, and (2) incorrect answer generation despite retrieving the correct context.
@@
The retriever returns the correct supporting triple spouse(Ian Holm, Penelope Wilton), which should allow the generator to infer the reverse direction. Nonetheless, the model produces the incorrect answer Marriage, likely due to interference from unrelated paths involving typeOfUnion relations. This suggests that even with accurate retrieval, models can fail to distinguish relevant from irrelevant context during generation.
@@ 6 Discussion and Conclusion
This work introduces a general methodology for constructing benchmarks aimed at evaluating the reasoning capabilities of KG-RAG systems under conditions of knowledge incompleteness—a realistic yet often overlooked challenge in real-world KGs. Our benchmark construction pipeline explicitly removes direct supporting facts while ensuring that answers remain inferable via alternative paths, allowing us to isolate and measure true reasoning performance.
Our empirical study on datasets derived from Family and FB15k-237 reveals several key limitations of current KG-RAG methods. Most notably, existing KG-RAG models struggle to recover answers when direct evidence is missing, indicating limited robustness to incomplete knowledge. While training-based methods exhibit greater resilience, our fine-grained analysis reveals potential overfitting to specific relation patterns, sometimes at the expense of generalizing over trivial structures. Furthermore, we find that textual entity labels substantially improve performance, suggesting that models rely more on retrieving memorized knowledge than performing symbolic reasoning over structured data.
These findings point to several promising directions for future research. First, there is a need for more advanced retrieval strategies that can effectively identify alternative paths when direct evidence is missing. Second, reasoning modules should be designed with greater generalization capability to avoid overfitting to specific relation patterns. Third, fine-tuning strategies should be carefully crafted to enhance retrieval without compromising the LLM’s inherent reasoning ability.
@@ 7 Limitation
One limitation of our benchmark lies in the expressiveness of the logical rules used for constructing reasoning questions. Specifically, we rely on rules mined by AMIE3, which are restricted to Horn-style rules. While this rule format is efficient and widely adopted, it captures only a subset of the reasoning patterns present in real-world KGs.
As a result, our benchmark primarily evaluates reasoning over relatively simple logical structures (e.g., symmetry, transitivity, composition). More sophisticated forms of inference—such as disjunctive reasoning, aggregation (e.g., counting), and numerical constraints—are not currently represented. These reasoning types are important in many real-world applications and pose greater challenges for KG-RAG systems.
In future work, it would be valuable to explore methods for incorporating richer rule types into the benchmark, either by extending the rule mining process (e.g., with neural-symbolic rule learners or probabilistic rule discovery techniques) or by manually curating complex inference templates. This would allow for a more comprehensive evaluation of KG-RAG models’ reasoning capabilities beyond basic logical chaining.
@@ A.1 AMIE3 Candidate Rule Refinement
AMIE3 generate candidate rules by a refinement process using a classical breadth-first search Lajus et al. (2020). It begins with rules that contain only a head atom (e.g. $\Rightarrow\texttt{hasSibling}(X,Y)$) and refines them by adding atoms to the body. For example, it may generate the refined rule:
@@ A.2 AMIE3 Hyperparameter Settings
We use AMIE3 with a confidence threshold of 0.3 and a PCA confidence threshold $\theta_{\text{PCA}}$ of 0.4 for both datasets. The maximum rule length is set to 3 for Family to avoid overly complex patterns, and 4 for FB15k-237 to allow richer rules. See Appendix B.1 for the definition of PCA confidence.
@@ Appendix B Properties of Horn Rules Mined by AMIE3
These restrictions are widely adopted in KG rule mining (Galárraga et al., 2015; Lajus et al., 2020) to guarantee that discovered rules are logically well-formed and meaningful for downstream reasoning tasks.
@@ B.1 PCA Confidence
Definition. To address this, AMIE3 introduces PCA confidence (Partial Completeness Assumption confidence) (Galárraga et al., 2015), an optimistic variant that partially compensates for KG incompleteness. Given a rule of the form
@@
For further details, see (Galárraga et al., 2015).
@@ B.2 Rule Mining Procedure
AMIE3 generate candidate rules by a refinement process using a classical breadth-first search Lajus et al. (2020). Algorithm 2 summarizes the rule mining process of AMIE3. The algorithm starts with an initial set of rules that contain only a head atom (i.e. $\top\Rightarrow r(X,Y)$, where $\top$ denotes an empty body) and maintains a queue of rule candidates (Line 1). At each step, AMIE3 dequeues a rule $R$ from the queue and evaluates whether it satisfies three criteria (Line 5):
@@ B.3 Benchmark Construction Code and Data
We release the source code for benchmark construction, along with the Family and FB15k-237 benchmark datasets, at https://anonymous.4open.science/r/INCK-EA16.
@@ Appendix C Additional Results of the experiment
Figure 5 presents the recall and precision of all evaluated KG-RAG models on the constructed benchmarks.
Figure 5: Performance comparison of KG-RAG models under incomplete (blue) and complete (red) KG settings, measured by Recall (top) and Precision (bottom).
@@ Appendix D Prompt Template
To ensure reproducibility and mitigate randomness in LLM outputs Potyka et al. (2024), we set the generation temperature to 0 in all experiments.
@@ Appendix F Baseline Details
RoG, G-Retriever, and GNN-RAG are each trained and evaluated separately on the 8:1:1 training split of each dataset (Family and FB15k-237) using a single NVIDIA H200 GPU, as described in Section 3.3. For RoG, we use LLaMA2-Chat-7B as the LLM backbone, instruction-finetuned on the training split of Family or FB15K-237 for 3 epochs. The batch size is set to 4, the learning rate to $2\times 10^{-5}$, and a cosine learning rate scheduler with a warmup ratio of 0.03 is adopted (Luo et al., 2024). For G-Retriever, the GNN backbone is a Graph Transformer (4 layers, 4 attention heads per layer, hidden size 1024) with LLaMA2-7B as the LLM. Retrieval hyperparameters and optimization follow He et al. (2024). For GNN-RAG (Mavromatis and Karypis, 2024), we use the recommended ReaRev backbone and sBERT encoder; the GNN component is trained for 200 epochs with 80 epochs of warmup and a patience of 5 for early stopping. All random seeds are fixed for reproducibility. For PoG (Chen et al., 2024), StructGPT (Jiang et al., 2023), and ToG (Sun et al., 2024), we use GPT-3.5-turbo as the underlying LLM, and the original prompt and generation settings from each method.
@@ Appendix G Detailed Analysis of Other Rule Types
Some rules do not exhibit simple interpretable motifs, involving unusual variable binding or rare predicate combinations. Like recursive rules (check AMIE3 Lajus et al. (2020) for more details)
Appendix H Personal Identification Issue in FB15k-237
While FB15k-237 contains information about individuals, it typically focuses on well-known public figures such as celebrities, politicians, and historical figures. Since this information is already widely available online and in various public sources, its inclusion in Freebase doesn’t significantly compromise individual privacy compared to datasets containing sensitive personal information.
Appendix I AI Assistants In Writing
We use ChatGPT OpenAI (2024) to enhance our writing skills, abstaining from its use in research and coding endeavors.

### v3

@@ 5.4 Influence of Entity Labeling
To assess how different entity labeling schemes affect KG-RAG performance, we evaluate three settings for representing entities in the input: (1) Private ID—randomly assigned IDs with no semantic content, (2) Entity ID—official Freebase IDs (e.g., /m/02mjmr for Barack Obama), and (3) Text Label—natural language labels of entities. Figure 4 shows the F1-scores of all models under each representation, for both incomplete (left) and complete (right) KG settings. We observe two key trends: First, text labels significantly boost performance. Models consistently achieve higher scores when entity labels are expressed in natural language. This suggests that LLMs can effectively leverage their internal knowledge when text labels are provided. Second, entity IDs provide limited benefit over random IDs. Surprisingly, using official entity IDs like /m/02mjmr results in performance nearly identical to that of randomly assigned private IDs. This indicates that, despite LLMs potentially memorizing mappings between surface forms and IDs, they are unable to reason with these identifiers in generation tasks. Instead, they appear to treat IDs as opaque tokens unless the text label is explicit.
Figure 4: F1-Score comparison of models using Private ID, Entity ID, and Text Label representations on FB15k-237 (a-b) and Wikidata5m (c-d) benchmark under incomplete and complete KG settings.
@@ 5.5 Case Study
To better understand the limitations of current KG-RAG methods, we analyze representative failure cases from our benchmark. Table 4 shows two illustrative examples: (1) failure to retrieve relevant reasoning paths and (2) incorrect answer generation despite correct retrieval. A quantitative analysis is provided in Appendix J.
@@
The retriever correctly retrieves spouse(Ian Holm, Penelope Wilton), which should enable inference of the reverse relation. However, the model outputs Marriage, likely influenced by unrelated typeOfUnion paths, indicating difficulty in filtering irrelevant context even with accurate retrieval.
@@ 6 Discussion and Conclusion
This work introduces a general methodology for constructing benchmarks aimed at evaluating KG-RAG systems under conditions of knowledge incompleteness—a realistic yet often overlooked challenge in real-world KGs.
Our empirical study on datasets derived from Family, FB15k-237 and Wikidata5m reveals several key limitations of current KG-RAG methods. Most notably, existing KG-RAG models struggle to recover answers when direct evidence is missing, indicating limited robustness to incomplete knowledge. While training-based methods exhibit greater resilience, our fine-grained analysis reveals potential overfitting to specific relation patterns, sometimes at the expense of generalizing over trivial structures. Furthermore, we find that textual entity labels substantially improve performance, suggesting that models rely more on retrieving memorized knowledge than performing symbolic reasoning over structured data.
These findings suggest several research directions: (1) developing retrieval strategies that identify alternative paths when evidence is missing, (2) designing reasoning modules with stronger generalization to avoid overfitting to specific relation patterns, and (3) fine-tuning methods that enhance retrieval without impairing reasoning.
@@ 7 Limitation
Our benchmark construction relies on Horn-style logical rules mined by AMIE3 to construct reasoning questions. While this format may appear restrictive, it in fact covers a broad range of reasoning patterns frequently observed in real-world KGs. As shown in Figure 3, the benchmark encompasses four major rule types that dominate real-world KGs, with additional diverse patterns summarized in Appendix G.
This focus on Horn rules also offers a practical advantage: such rules are interpretable and easy to verify for plausibility, ensuring the reliability and consistency of the constructed questions. In contrast, more expressive rule types are challenging to mine reliably, and even SOTA rule mining methods often produce noisy or implausible rules. To maintain benchmark quality, we therefore focus on Horn rules in this paper. Extending to richer rule types is a promising direction for future work, specifically, incorporating advanced rule-learning methods like DRUM (sadeghian2019drum) and AnyBURL (meilicke2024anytime), or aggregating high-confidence rules from multiple algorithms, offers a principled way to expand coverage beyond the capabilities of any single system.
@@ A.1 AMIE3 Candidate Rule Refinement
AMIE3 generates candidate rules by a refinement process using a classical breadth-first search lajus2020fast. It begins with rules that contain only a head atom (e.g. $\Rightarrow\texttt{hasSibling}(X,Y)$) and refines them by adding atoms to the body. For example, it may generate the refined rule:
@@ A.2 AMIE3 Hyperparameter Settings
We use AMIE3 with a confidence threshold of 0.3 and a PCA confidence threshold $\theta_{\text{PCA}}$ of 0.4 for all three datasets. The maximum rule length is set to 3 for Family to avoid overly complex patterns, and 4 for FB15k-237 and Wikidata5m to allow richer rules. For the AMIE3 confidence and head coverage thresholds (Section 3.1), we tested values from 0.1–0.6 and chose 0.3 as a balance: lower values produced unreliable rules, while higher ones biased the distribution toward certain relations. For $\tau$, we tested 0.01–0.2 and set it to limit domination by frequent entities, ensuring a balanced answer distribution. See Appendix B.1 for the definition of PCA confidence.
@@ Appendix B Properties of Horn Rules Mined by AMIE3
These restrictions are widely adopted in KG rule mining (galarraga2015fast; lajus2020fast) to guarantee that discovered rules are logically well-formed and meaningful for downstream reasoning tasks.
@@ B.1 PCA Confidence
Definition. To address this, AMIE3 introduces PCA confidence (Partial Completeness Assumption confidence) (galarraga2015fast), an optimistic variant that partially compensates for KG incompleteness. Given a rule of the form
@@
For further details, see (galarraga2015fast).
@@ B.2 Rule Mining Procedure
AMIE3 generates candidate rules by a refinement process using a classical breadth-first search lajus2020fast. Algorithm 2 summarizes the rule mining process of AMIE3. The algorithm starts with an initial set of rules that contain only a head atom (i.e. $\top\Rightarrow r(X,Y)$, where $\top$ denotes an empty body) and maintains a queue of rule candidates (Line 1). At each step, AMIE3 dequeues a rule $R$ from the queue and evaluates whether it satisfies three criteria (Line 5):
@@ B.3 Benchmark Construction Code and Data
We release the source code for benchmark construction, along with the Family, FB15k-237 and Wikidata5m benchmark datasets, at https://anonymous.4open.science/r/INCK-EA16.
@@ Appendix C Additional Results of the experiment
C.1 Recall and Precision
In complement to Figure 2, we provide a detailed breakdown of recall and precision for all settings to offer a more comprehensive evaluation of the F1 scores. Figure 5 presents the recall and precision values for all evaluated KG-RAG models across the constructed benchmarks.
Figure 5: Performance comparison of KG-RAG models under incomplete (blue) and complete (red) KG settings, measured by Recall (top) and Precision (bottom) on Family, FB15k-237, and Wikidata5m.
C.2 Evaluation of Generate-on-Graph (GoG) (xu2024generate)
This section provides a detailed analysis of GoG (xu2024generate), following the experimental results presented in the main text. While GoG explicitly addresses knowledge incompleteness as part of its motivation, we did not include this aspect in the main body for several reasons.
GoG’s implementation is heavily optimized for specific datasets, such as CWQ (Talmor2018WSQ) and WebQSP (Yih2016webqsp), and relies on several domain-specific heuristics:
  • Hard-coded Templates: Prompting mechanisms and relation-name heuristics tailored to specific KG schemas.
  • Entity-Type Linking: Reliance on entity-type information that is often unavailable in general-purpose or incomplete KGs.
  • Textual Dependency: Assumptions regarding entity text labels, which prevents the model from functioning on datasets using anonymized numeric IDs—a necessary measure we take to prevent knowledge leakage and evaluate pure structural reasoning.
Despite these limitations, we adapted GoG for our benchmark. The results are shown in Table 5. Our evaluation indicates a significant performance decline, particularly in HHR, our metric designed to assess reasoning capacity when direct supporting facts are missing.
The primary limitation lies in GoG’s underlying reasoning mechanism. Rather than performing multi-hop reasoning over alternative paths within the KG, GoG tends to compensate for missing triples by querying the LLM’s internal parametric priors. While this strategy may succeed in benchmarks with significant knowledge overlap between the KG and the LLM’s training data, it fails in our setting. By anonymizing entity names, we isolate the model’s ability to reason over the provided structure.
Table 5: Performance of GoG on Family and FB15k-237.
Dataset Hits@Any HHR F1
Family Complete 0.60 0.96 0.48
Family Incomplete 0.58 0.28 0.36
FB15k-237 Complete 0.46 0.57 0.26
FB15k-237 Incomplete 0.20 0.23 0.11
@@ Appendix D Prompt Template
To ensure reproducibility and mitigate randomness in LLM outputs potyka2024robust, we set the generation temperature to 0 in all experiments.
@@ Appendix F Baseline Details
RoG, G-Retriever, and GNN-RAG are each trained and evaluated separately on the 8:1:1 training split of each dataset (Family, FB15k-237 and Wikidata5m) using a single NVIDIA H200 GPU, as described in Section 3.3. For RoG, we use LLaMA2-Chat-7B as the LLM backbone, instruction-finetuned on each dataset for 3 epochs separately. The batch size is set to 4, the learning rate to $2\times 10^{-5}$, and a cosine learning rate scheduler with a warmup ratio of 0.03 is adopted (linhao2024rog). For G-Retriever, the GNN backbone is a Graph Transformer (4 layers, 4 attention heads per layer, hidden size 1024) with LLaMA2-7B as the LLM. Retrieval hyperparameters and optimization follow he2024g. For GNN-RAG (mavromatis2024gnn), we use the recommended ReaRev backbone and sBERT encoder; the GNN component is trained for 200 epochs with 80 epochs of warmup and a patience of 5 for early stopping. All random seeds are fixed for reproducibility. For PoG (chen2024pog), StructGPT (jiang2023structgpt), and ToG (Jiashuo2024ToG), we use GPT-3.5-turbo as the underlying LLM, following the original papers, and adopt the original prompt and generation settings from each method. The temperature of LLMs is set to 0.
@@ Appendix G Detailed Analysis of Other Rule Types
Some rules do not exhibit simple interpretable motifs, involving unusual variable binding or rare predicate combinations. Like recursive rules (check AMIE3 lajus2020fast for more details)
Appendix H Quality Check of the Generated Questions
To ensure that the constructed benchmark does not produce unanswerable questions, we conduct a quality check on the generated question set. Although the construction process guarantees that all grounded body triples required to infer a removed head remain in the KG, it is still necessary to verify that the removed triples can indeed be inferred from the remaining facts. Ensuring the answerability of generated questions is therefore essential, and a quantitative verification is required to confirm that the removed triples remain inferable from the retained facts.
To examine this, we evaluate whether each removed triple can be recovered when the model is given sufficient contextual information. Specifically, we provide both the alternative reasoning paths and the corresponding logical rules for each question, thereby simulating a perfect retriever with guided reasoning. This setup ensures that all necessary evidence for deriving the target entity is explicitly available. Performance is reported as the proportion of cases in which the target entity was correctly predicted.
The results in Table 6 show that, under this setting, LLMs answer most questions correctly, indicating that the benchmark questions are indeed answerable by LLMs when sufficient information is available.
Table 6: Performance of backbone LLMs when provided with sufficient information
Backbone Family FB15k-237 Wikidata5m
GPT-3.5-turbo 0.88 0.86 0.84
GPT-4o 0.91 0.94 0.90
Appendix I Results under Permissive Evaluation Metrics
We also report the performance of representative KG-RAG models under the commonly used permissive metrics adopted in prior works, such as Hits@1 and relaxed F1/Recall/Precision (linhao2024rog; chen2024pog). These metrics treat partial string matches as correct answers and may therefore overestimate model performance, for example by counting a response that contains the gold entity in a negated form as correct (e.g., treating “not Barack Obama” as correct when the gold answer is “Barack Obama”). Nevertheless, as shown in Table 7, the results remain consistent with those obtained using our stricter evaluation protocol, revealing substantial performance differences between complete and incomplete KGs.
Table 7: Performance of representative KG-RAG models under standard permissive metrics.
Method Settings Hits@1 Permissive F1 Permissive Recall Permissive Precision
RoG complete KG 69.54 43.97 47.25 46.08
RoG incomplete KG 62.97 33.24 33.63 37.81
PoG complete KG 57.59 30.21 27.49 56.99
PoG incomplete KG 45.26 18.56 18.47 45.66
The consistent performance gap across both permissive and strict evaluation settings further validates the robustness of our findings and highlights the distinct behaviors of KG-RAG models when reasoning over incomplete knowledge graphs. In particular, both representative methods (RoG and PoG) exhibit significant performance drops across permissive and strict evaluation settings, and other models show similarly clear declines under incomplete KGs.
Appendix J Quantitative Error Analysis
To complement the qualitative case studies presented in Section 5.5, we further quantify the distribution of failure types across datasets and methods. Based on the taxonomy introduced in Section 5.5, we aggregate the results into two broad categories: Retrieval Failure and Reasoning Failure. On average, the ratio between these two categories is approximately 7:3, with the majority of failures attributable to Retrieval Failure rather than Reasoning Failure. The ratio is averaged across datasets and models, following the classification criteria in Section 5.5: a case is marked as Retrieval Failure if no alternative path supporting the gold answer is retrieved, and as Reasoning Failure if such a path is retrieved but the final answer is incorrect.
This quantitative summary reinforces the qualitative findings discussed in Section 5.5, suggesting that Retrieval Failure remains the primary bottleneck for KG-RAG models operating under incomplete KGs.
Appendix K Personal Identification Issue in FB15k-237 and Wikidata5m
While FB15k-237 and Wikidata5m contain information about individuals, they typically focus on well-known public figures such as celebrities, politicians, and historical figures. Since this information is already widely available online and in various public sources, their inclusion in Freebase and Wikidata doesn’t significantly compromise individual privacy compared to datasets containing sensitive personal information.
Appendix L AI Assistants In Writing
We use ChatGPT openai2024chatgpt to enhance our writing skills, abstaining from its use in research and coding endeavors.